# Preasymptotic error estimates of higher-order EEM for the time-harmonic Maxwell equations with large wave number

### v1

@@ 2.2. The EEM
The well-posedness of the above problem is guaranteed by [46, Theorem 4.17].
@@
[…]
@@
Here, $\mathcal{P}_{p}(\hat{K})$ denotes the set of polynomials of degree less than or equal to $p$ on $\hat{K}$. The EEM for the Maxwell problem (1.1) reads as: Find ${\bm{E}}_{h}\in\bm{V}_{h}$ such that
@@ 2.4. Helmholtz decompositions and embedding results
The following well-known decomposition of ${\bm{L}}^{2}{(\Omega)}$ will be frequently used in our subsequent analyses, the proof of which can be found in [6, Proposition 3.7.3] or [53, Theorem 4.2].
@@
Lemma 2.4 (Helmholtz decomposition of ${\bm{V}_{h}^{0}}$).
@@
For the proof of this lemma, we refer to [46, Theorem 3.45] and [46, Lemma 7.6].
@@ 2.5. Stability and regularity results
$\|\bm{E}\|_{l}\lesssim\kappa^{l-1}C_{l-1,\bm{f},\bm{g}},\quad 2\leq l\leq m-1.$
Recall that $\operatorname{{\mathrm{div}}}_{\Gamma}(\bm{v}\times\bm{\nu})=\bm{\nu}\cdot\operatorname{\bm{\mathrm{curl}}}\bm{v}$ (see, e.g., [46, (3.52)]). We have from (1.1b) and (1.1a) that
@@
(a) The regularity estimate (2.19) was proved in [36, Theorem 3.4] when the domain $\Omega$ is $C^{m+1}$ and in [43, Lemma 5.1] when $\Omega$ is sufficiently smooth. [19, Theorem 1] gave a regularity estimate by assuming only $C^{m}$-domain, however, explicit dependence on the wave number $\kappa$ was not considered there.
@@
3. An auxiliary Maxwell elliptic problem
In this section, we first construct a truncation operator for any divergence-free function by truncating its $\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}$ eigenfunction expansion. Using this operator, we define an auxiliary problem and present key regularity results that are essential for the subsequent preasymptotic analysis.
@@ 3.1. A truncation operator
Let $\mathcal{P}:=\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}-\nabla\operatorname{{\mathrm{div}}}$. It is easy verify that ${\bm{\phi}}_{j}$ is the unique weak solution in ${\bm{X}}$ of the following boundary value problem:
$\mathcal{P}{\bm{\phi}}&=\lambda_{j}{\bm{\phi}}_{j}&\text{ in }\Omega,\\ \operatorname{\bm{\mathrm{curl}}}{\bm{\phi}}\times\bm{\nu}&=\bm{0}&\text{ on }\Gamma,\\ {\bm{\phi}}\cdot\bm{\nu}&=\bm{0}&\text{ on }\Gamma.\end{aligned}\right.$ (3.2)
@@
for $\text{for any }\bm{v}=\sum_{j=1}^{\infty}\bm{v}_{j}{\bm{\phi}}_{j}\in\bm{H}(\operatorname{{\mathrm{div}}}^{0};\Omega),\quad\text{let }\mathcal{T}\bm{v}=\sum_{j=1}^{N}\bm{v}_{j}{\bm{\phi}}_{j},$
@@
For any $m\in\mathbb{N}_{1}$, suppose $\Omega$ is $C^{\max\{2,m\}}$. Then we have
@@ 3.2. The auxiliary Maxwell elliptic problem
Using the fact that $\operatorname{\bm{\mathrm{curl}}}{\bm{u}}_{1}={\bm{0}}$ and the Young’s inequality, we have
@@
It is straightforward to observe that the auxiliary problem mentioned above differs from (1.1) by adding some smoothed, divergence-free component to the left-hand side of the original equation. Compared to the original problem (see Theorem 2.1), the solution to the auxiliary problem exhibits improved behavior with respect to $\kappa$-dependence. Specifically, we have the following lemma.
@@
$\|{\bm{w}}\|_{1}$ $\lesssim\kappa^{-1}\|{\bm{f}^{\prime}}\|+\kappa^{-\frac{1}{2}}\|{\bm{g}^{\prime}}\|_{\Gamma}+\kappa^{-1}\|\operatorname{{\mathrm{div}}}{\bm{g}^{\prime}}\|_{-\frac{1}{2},\Gamma}.$ (3.14)
@@
$\tilde{C}_{m-2,{\bm{f}^{\prime}},{\bm{g}^{\prime}}}$ $:=\kappa^{-m+2}\big{(}\|{\bm{f}^{\prime}}\|_{m-2,\kappa}+\|{\bm{g}^{\prime}}\|_{m-\frac{3}{2},\kappa}\big{)}+\kappa^{\frac{1}{2}}\|{\bm{g}^{\prime}}\|_{\Gamma}+\kappa^{-m+1}\big{(}\|{\bm{f}^{\prime}}\cdot\bm{\nu}\|_{m-\frac{3}{2},\Gamma}+\|\operatorname{{\mathrm{div}}}{\bm{g}^{\prime}}\|_{m-\frac{3}{2},\Gamma}\big{)}.$
The stability result in (3.13) follows directly from (3.9), while (3.14)–(3.16) can be established using (3.13) along with techniques employed for demonstrating the regularity of the original Maxwell problem. For the reader’s convenience, the detailed proof of Lemma 3.3 is provided in Appendix A.
@@
(a) Let $\Omega\in C^{m}$ for some integer $m\geq 2$, if ${\bm{f}^{\prime}}\in{\bm{H}}^{m-2}(\operatorname{{\mathrm{div}}}^{0};\Omega)$, ${\bm{f}^{\prime}}\cdot\bm{\nu}=0$ on $\Gamma$, and ${\bm{g}^{\prime}}=\bm{0}$, then we have $\|{\bm{w}}\|_{m}\lesssim\|{\bm{f}^{\prime}}\|_{m-2,\kappa}$.
@@
(b) Similar to the original continuous problem, if $\operatorname{{\mathrm{div}}}{\bm{f}^{\prime}}\neq 0$, the stability estimate (3.13) still holds, the regularity estimate (3.14) and (3.15) hold if an additional term $\kappa^{-1}\|\operatorname{{\mathrm{div}}}{\bm{f}^{\prime}}\|$ and $\|\operatorname{{\mathrm{div}}}{\bm{f}^{\prime}}\|$ is added to the right-hand side of (3.14) and (3.15), respectively. The ${\bm{H}}^{m}$ regularity estimate (3.16) holds if the domain $\Omega$ is $C^{m+1}$ and an additional term $\kappa^{-m}\big{(}\|\operatorname{{\mathrm{div}}}{\bm{f}^{\prime}}\|_{m-1}+\kappa\|\operatorname{{\mathrm{div}}}{\bm{f}^{\prime}}\|_{m-2}\big{)}$ is added to $\tilde{C}_{m-2,{\bm{f}^{\prime}},{\bm{g}^{\prime}}}$.
4. Preasymptotic error estimates for the EEM
In this section, we first establish some wave-number-explicit stability estimates for the $p$-th order EEM (2.5) by using the so-called “modified duality argument” developed for the preasymptotic analysis of FEM for Helmholtz equation [57]. Then we use the resulting stability estimates to derive preasymptotic error estimates for the EEM. The proof is generally in the same spirit of its linear counterpart (see [38, Theorem 4.1 and 4.2]), but differs in the projection operators $\bm{P}^{\pm}_{h}$.
In the “modified duality argument” step, the following Maxwell dual problem and the $\kappa$-explicit approximation (see Lemma 4.1) of its solution plays a vital role.
4.1. $\kappa$-explicit regularity by decomposition for the Maxwell dual problem
We introduce the following Maxwell dual problem:
$\left\{\begin{array}[]{rlrl}\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}{\bm{\Psi}}-\kappa^{2}{\bm{\Psi}}&={\bm{f}^{\prime}}&\text{ in}\Omega,\\ \operatorname{\bm{\mathrm{curl}}}{\bm{\Psi}}\times\bm{\nu}+\bm{\mathrm{i}}\kappa\lambda{\bm{\Psi}}_{T}&=\bm{0}&\text{ on }\Gamma,\end{array}\right.$ (4.1)
where ${\bm{f}^{\prime}}\in\bm{H}(\operatorname{{\mathrm{div}}}^{0};\Omega)$.
The following theorem says that the solution ${\bm{\Psi}}$ to the Maxwell dual problem (4.1) can be decomposed into the sum of an elliptic part and an analytic part: ${\bm{\Psi}}={{\bm{\Psi}}}_{\mathcal{E}}+{{\bm{\Psi}}}_{\mathcal{A}}$, where ${{\bm{\Psi}}}_{\mathcal{E}}$ is usually nonsmooth but the ${\bm{H}}^{2}$-norm of ${{\bm{\Psi}}}_{\mathcal{E}}$ is uniformly bounded for all ${\kappa}$ and ${{\bm{\Psi}}}_{\mathcal{A}}$ is oscillatory but the ${\bm{H}}^{j}$-bound of ${{\bm{\Psi}}}_{\mathcal{A}}$ is available for any integer $j\geq 0$ when the domain is sufficiently smooth.
Theorem 4.1 (Regularity decomposition).
Suppose that $\Omega$ is $C^{m}$ for some integer $m\geq 2$ and ${\bm{f}^{\prime}}\in{\bm{H}}(\operatorname{{\mathrm{div}}}^{0};\Omega)$. Let ${\bm{\Psi}}$ be the solution to problem (4.1), then there exists a splitting ${\bm{\Psi}}={{\bm{\Psi}}}_{\mathcal{E}}+{{\bm{\Psi}}}_{\mathcal{A}}$ such that ${{\bm{\Psi}}}_{\mathcal{E}}\in{\bm{H}}^{1}(\Omega)$ and ${{\bm{\Psi}}}_{\mathcal{A}}\in{\bm{H}}^{m}(\Omega)$ satisfying
$\kappa^{1-m}\|{{\bm{\Psi}}}_{\mathcal{A}}\|_{m}+\kappa^{2}\|{{\bm{\Psi}}}_{\mathcal{E}}\|+\kappa\|{{\bm{\Psi}}}_{\mathcal{E}}\|_{1}\lesssim\|{\bm{f}^{\prime}}\|.$ (4.2)
Furthermore, if ${\bm{f}^{\prime}}\cdot\bm{\nu}\in{\bm{H}}^{\frac{1}{2}}(\Gamma)$, then ${{\bm{\Psi}}}_{\mathcal{E}}\in{\bm{H}}^{2}(\Omega)$ and
$\|{{\bm{\Psi}}}_{\mathcal{E}}\|_{2}\lesssim\|{\bm{f}^{\prime}}\|+\kappa^{-1}\|{\bm{f}^{\prime}}\cdot\bm{\nu}\|_{\frac{1}{2},\Gamma}.$ (4.3)
Proof.
Let ${{\bm{\Psi}}}_{\mathcal{E}}$ be the solution to the following problem:
$\begin{cases}\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}{{\bm{\Psi}}}_{\mathcal{E}}-\kappa^{2}{{\bm{\Psi}}}_{\mathcal{E}}+C_{\mathcal{T}}\kappa^{2}\mathcal{T}\textsf{P}^{\perp}{{\bm{\Psi}}}_{\mathcal{E}}={\bm{f}^{\prime}}&\text{ in }\Omega,\\ \operatorname{\bm{\mathrm{curl}}}{{\bm{\Psi}}}_{\mathcal{E}}\times\bm{\nu}+\bm{\mathrm{i}}\kappa\lambda({{\bm{\Psi}}}_{\mathcal{E}})_{T}=\bm{0}&\text{ on }\Gamma.\end{cases}$ (4.4)
The proof of Lemma 3.3 shows that the estimate for ${{\bm{\Psi}}}_{\mathcal{E}}$ in (4.2)–(4.3) holds.
Let ${{\bm{\Psi}}}_{\mathcal{A}}={\bm{\Psi}}-{{\bm{\Psi}}}_{\mathcal{E}}$. It is obvious that ${{\bm{\Psi}}}_{\mathcal{A}}$ satisfies
$\begin{cases}\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}{{\bm{\Psi}}}_{\mathcal{A}}-\kappa^{2}{{\bm{\Psi}}}_{\mathcal{A}}=C_{\mathcal{T}}\kappa^{2}\mathcal{T}\textsf{P}^{\perp}{{\bm{\Psi}}}_{\mathcal{E}}&\text{ in }\Omega,\\ \operatorname{\bm{\mathrm{curl}}}{{\bm{\Psi}}}_{\mathcal{A}}\times\bm{\nu}+\bm{\mathrm{i}}\kappa\lambda({{\bm{\Psi}}}_{\mathcal{A}})_{T}=\bm{0}&\text{ on }\Gamma.\end{cases}$ (4.5)
By the regularity results in Theorem 2.1 and the stability of $\mathcal{T}$ and $\textsf{P}^{\perp}$, we have
$\kappa^{-1}\|{{\bm{\Psi}}}_{\mathcal{A}}\|_{1}+\|{{\bm{\Psi}}}_{\mathcal{A}}\|$ $\lesssim C_{\mathcal{T}}\kappa\|\mathcal{T}\textsf{P}^{\perp}{{\bm{\Psi}}}_{\mathcal{E}}\|\lesssim\kappa^{-1}\|{\bm{f}^{\prime}}\|,$ (4.6)
$\|{{\bm{\Psi}}}_{\mathcal{A}}\|_{m}$ $\lesssim C_{\mathcal{T}}\kappa^{m+1}\big{(}\|\mathcal{T}\textsf{P}^{\perp}{{\bm{\Psi}}}_{\mathcal{E}}\|+\sum_{i=1}^{m-2}\kappa^{-1-i}\|\mathcal{T}\textsf{P}^{\perp}{{\bm{\Psi}}}_{\mathcal{E}}\|_{i}+\kappa^{-m}\|(\mathcal{T}\textsf{P}^{\perp}{{\bm{\Psi}}}_{\mathcal{E}})\cdot\bm{\nu}\|_{m-\frac{3}{2},\Gamma}\big{)}$
$\lesssim\kappa^{m-1}\|{\bm{f}^{\prime}}\|.$ (4.7)
Then the proof is completed. ∎
Remark 4.1.
(a) As discussed in [24, 34], the above proof presents a more straightforward methodology compared to existing regularity splitting approaches, such as the Fourier technique [40, 41, 42, 43] and iterative techniques found in [16, 17].
The following lemma provides some approximation estimates for the solution to the Maxwell dual problem (4.1), relying on its regularity decomposition.
Lemma 4.1.
Suppose that $\Omega\in C^{p+1}$, ${\bm{f}^{\prime}}\in{\bm{H}}(\operatorname{{\mathrm{div}}}^{0};\Omega)$ and ${\bm{f}^{\prime}}\cdot\bm{\nu}\in{\bm{H}}^{\frac{1}{2}}(\Gamma)$. Let ${\bm{\Psi}}$ be the solution to (4.1), then there exists $\widehat{{\bm{\Psi}}}_{h}\in\bm{V}_{h}$ such that if $\kappa h\lesssim 1$, then
$h^{\frac{1}{2}}\|({\bm{\Psi}}-\widehat{{\bm{\Psi}}}_{h})_{T}\|_{\Gamma}+\|{\bm{\Psi}}-\widehat{{\bm{\Psi}}}_{h}\|+h\interleave{\bm{\Psi}}-\widehat{{\bm{\Psi}}}_{h}\interleave\lesssim h^{2}\big{(}\|{\bm{f}^{\prime}}\|+\kappa^{-1}\|{\bm{f}^{\prime}}\cdot\bm{\nu}\|_{\frac{1}{2},\Gamma}\big{)}+\kappa^{p}h^{p+1}\|{\bm{f}^{\prime}}\|.$ (4.8)
Proof.
By using Theorem 4.1, there exists a splitting ${\bm{\Psi}}={{\bm{\Psi}}}_{\mathcal{E}}+{{\bm{\Psi}}}_{\mathcal{A}}$ satisfying
$\|{{\bm{\Psi}}}_{\mathcal{E}}\|_{2}\lesssim\|{\bm{f}^{\prime}}\|+\kappa^{-1}\|{\bm{f}^{\prime}}\cdot\bm{\nu}\|_{\frac{1}{2},\Gamma},$
$\|{{\bm{\Psi}}}_{\mathcal{A}}\|_{p+1}\lesssim\kappa^{p}\|{\bm{f}^{\prime}}\|.$
Then by the Lemma 2.1 and using the fact that $\kappa h\lesssim 1$, there exist ${\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{E}}$ and ${\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{A}}\in\bm{V}_{h}$ such that
$h^{\frac{1}{2}}\|({{\bm{\Psi}}}_{\mathcal{E}}-{\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{E}})_{T}\|_{\Gamma}+\|{{\bm{\Psi}}}_{\mathcal{E}}-{\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{E}}\|$ $\lesssim h^{2}\|{{\bm{\Psi}}}_{\mathcal{E}}\|_{2},$
$h(\|\operatorname{\bm{\mathrm{curl}}}({{\bm{\Psi}}}_{\mathcal{E}}-{\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{E}})\|+\kappa^{\frac{1}{2}}\|({{\bm{\Psi}}}_{\mathcal{E}}-{\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{E}})_{T}\|_{\Gamma})$ $\lesssim h^{2}\|{{\bm{\Psi}}}_{\mathcal{E}}\|_{2},$
$h^{\frac{1}{2}}\|({{\bm{\Psi}}}_{\mathcal{A}}-{\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{A}})_{T}\|_{\Gamma}+\|{{\bm{\Psi}}}_{\mathcal{A}}-{\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{A}}\|$ $\lesssim h^{p+1}\|{{\bm{\Psi}}}_{\mathcal{A}}\|_{p+1},$
$h(\|\operatorname{\bm{\mathrm{curl}}}({{\bm{\Psi}}}_{\mathcal{A}}-{\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{A}})\|+\kappa^{\frac{1}{2}}\|({{\bm{\Psi}}}_{\mathcal{A}}-{\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{A}})_{T}\|_{\Gamma})$ $\lesssim h^{p+1}\|{{\bm{\Psi}}}_{\mathcal{A}}\|_{p+1}.$
Setting $\widehat{{\bm{\Psi}}}_{h}:={\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{E}}+{\bm{\Pi}_{h}}{\bm{\Psi}}_{\mathcal{A}}\in\bm{V}_{h}$ and using the triangle inequality, we obtain (4.8). ∎

### v2

@@ 2.2. The EEM
The well-posedness of the above problem is guaranteed by [47, Theorem 4.17].
@@
$h_{K}^{-|\alpha|}\left\|\mathrm{D}^{\alpha}F_{K}\right\|_{L^{\infty}(\hat{K})}\leq C\quad\forall K\in\mathcal{M}_{h},\;1\leq\left|\alpha\right|\leq p+2,$
@@
[…]
@@ 2.4. Helmholtz decompositions and embedding results
The following well-known decomposition of ${\bm{L}}^{2}{(\Omega)}$ will be frequently used in our subsequent analyses, the proof of which can be found in [6, Proposition 3.7.3] or [54, Theorem 4.2].
@@
Lemma 2.4 (Helmholtz decompositions of $\bm{V}_{h}^{0}$}).
@@
For the proof of this lemma, we refer to [47, Theorem 3.45] and [47, Lemma 7.6].
@@ 2.5. Stability and regularity results
$\|\bm{E}\|_{l}\lesssim\kappa^{l-1}C_{l-2,\bm{f},\bm{g}},\quad 2\leq l\leq m-1.$
Recall that $\operatorname{{\mathrm{div}}}_{\Gamma}(\bm{v}\times\bm{\nu})=\bm{\nu}\cdot\operatorname{\bm{\mathrm{curl}}}\bm{v}$ (see, e.g., [47, (3.52)]). We have from (1.1b) and (1.1a) that
@@
(a) The regularity estimate (2.19) was proved in [36, Theorem 3.4] (and in [44] for heterogeneous media) when the domain $\Omega$ is $C^{m+1}$ and in [43, Lemma 5.1] when $\Omega$ is sufficiently smooth. [19, Theorem 1] gave a regularity estimate by assuming only $C^{m}$-domain, however, explicit dependence on the wave number $\kappa$ was not considered there.
@@
3. An auxiliary Maxwell elliptic problem and elliptic projections
In this section, we first construct a truncation operator for any divergence-free function by truncating its $\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}$ eigenfunction expansion. Then using this operator, we define an auxiliary problem and present key regularity estimates. Finally, we introduce corresponding elliptic projections and derive their error estimates. The results obtained in this section are essential for the subsequent preasymptotic analysis.
@@ 3.1. A truncation operator
Let $\mathcal{P}:=\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}-\nabla\operatorname{{\mathrm{div}}}$. It is easy to verify that ${\bm{\phi}}_{j}$ is the unique weak solution in $\bm{X}$} of the following boundary value problem:
$\left\{\begin{aligned} \mathcal{P}{\bm{\phi}}&=\lambda_{j}{\bm{\phi}}_{j}&\text{ in }\Omega,\\ \operatorname{\bm{\mathrm{curl}}}{\bm{\phi}}\times\bm{\nu}&=\bm{0}&\text{ on }\Gamma,\\ {\bm{\phi}}\cdot\bm{\nu}&=\bm{0}&\text{ on }\Gamma.\end{aligned}\right$. (3.2)
@@
$\text{for any }\bm{v}=\sum_{j=1}^{\infty}\bm{v}_{j}{\bm{\phi}}_{j}\in\bm{H}(\operatorname{{\mathrm{div}}}^{0};\Omega),\quad\text{let }\mathcal{T}\bm{v}=\sum_{j=1}^{N}\bm{v}_{j}{\bm{\phi}}_{j},$
@@
Given $m\in\mathbb{N}_{1}$, suppose $\Omega$ is $C^{\max\{2,m\}}$. Then we have
@@ 3.2. The auxiliary Maxwell elliptic problem
Using the fact that $\operatorname{\bm{\mathrm{curl}}}{\bm{u}}_{1}={\bm{0}}$, $({\bm{u}}_{1},{\bm{u}}_{2})=0$, and the Young’s inequality, we have
@@
It is straightforward to observe that the auxiliary problem mentioned above differs from (1.1) by adding some smoothed, divergence-free component to the left-hand side of the original equation. Compared to the original problem (see Theorem 2.1), the solution to the auxiliary problem exhibits improved behavior with respect to $\kappa$-dependence. Specifically, we have the following lemma whose proof can be obtained by using standard techniques for regularity estimates for Maxwell equations (see e.g. [27, 36]) and is postponed to Appendix A.
@@
$\|{\bm{w}}\|_{1}$ $\lesssim\kappa^{-1}\|{\bm{f}^{\prime}}\|+\kappa^{-\frac{1}{2}}\|{\bm{g}^{\prime}}\|_{\Gamma}+\kappa^{-1}\|{\operatorname{{\mathrm{div}}}_{\Gamma}}{\bm{g}^{\prime}}\|_{-\frac{1}{2},\Gamma}.$ (3.14)
@@
$\tilde{C}_{m-2,{\bm{f}^{\prime}},{\bm{g}^{\prime}}}$ $:=\kappa^{-m+2}\big(\|{\bm{f}^{\prime}}\|_{m-2,\kappa}+\|{\bm{g}^{\prime}}\|_{m-\frac{3}{2},\kappa{,\Gamma}}\big)$
$\qquad+\kappa^{\frac{1}{2}}\|{\bm{g}^{\prime}}\|_{\Gamma}+\kappa^{-m+1}\big(\|{\bm{f}^{\prime}}\cdot\bm{\nu}\|_{m-\frac{3}{2},\Gamma}+\|{\operatorname{{\mathrm{div}}}_{\Gamma}}{\bm{g}^{\prime}}\|_{m-\frac{3}{2},\Gamma}\big).$
@@
(a) If $\Omega\in C^{m}$ for some integer $m\geq 2$, ${\bm{f}^{\prime}}\in{\bm{H}}^{m-2}(\operatorname{{\mathrm{div}}}^{0};\Omega)$, ${\bm{f}^{\prime}}\cdot\bm{\nu}=0$ and ${\bm{g}^{\prime}}=\bm{0}$ on $\Gamma$, then we have $\|{\bm{w}}\|_{m}\lesssim\|{\bm{f}^{\prime}}\|_{m-2,\kappa}$.
@@
(b) Similar to the original continuous problem, if $\operatorname{{\mathrm{div}}}{\bm{f}^{\prime}}\neq 0$, the stability estimate (3.13) still holds, the regularity estimate (3.14) and (3.15) hold if an additional term $\kappa^{-2}\|\operatorname{{\mathrm{div}}}{\bm{f}^{\prime}}\|$ and $\kappa^{-1}\|\operatorname{{\mathrm{div}}}{\bm{f}^{\prime}}\|$ is added to the right-hand side of (3.14) and (3.15), respectively. The ${\bm{H}}^{m}$ regularity estimate (3.16) holds if the domain $\Omega$ is $C^{m+1}$ and an additional term $\kappa^{-m}\big(\|\operatorname{{\mathrm{div}}}{\bm{f}^{\prime}}\|_{m-1}+\kappa\|\operatorname{{\mathrm{div}}}{\bm{f}^{\prime}}\|_{m-2}\big)$ is added to $\tilde{C}_{m-2,{\bm{f}^{\prime}},{\bm{g}^{\prime}}}$.
3.3. Elliptic projections